y=3x
x+y=4500 solve this by elimination

Answer by stanbon(75887)

y=3x
x+y=4500 solve this by elimination
----
Rearrange:
3x-y = 0
x +y = 4500
-----
Add ans solve for "x":
4x = 4500
x = 1125
-----
Solve for "y":
y = 3x
y = 3*1125 = 3375
=====================
Cheers,
Stan H.
========